OpenStudy (geekfromthefutur):

The table below shows the values of f(n) for different values of n: n 1 2 3 4 5 6 f(n) 1 2 4 16 128 4096 Which recursive function best represents the values shown in the table? f(1) = 1, f(2) = 2, f(n) = f(n -3) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = 2f(n -1) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = 2f(n -1) f(n - 2); n > 2

OpenStudy (geekfromthefutur):

@UnkleRhaukus

OpenStudy (unklerhaukus):

powers of two right?

OpenStudy (geekfromthefutur):

yea (:

OpenStudy (unklerhaukus):

we can't know f(n-3) for n=3 because we dont have f(0)

OpenStudy (unklerhaukus):

lets check that last option f(1) = 1, f(2) = 2, f(3) = 2f(3 -1) f(3 - 2) = ? does this equal 4? f(4) = 2f(4 -1) f(4 - 2) = ? does this equal 16?

OpenStudy (geekfromthefutur):

the first one no it does not aand the second one does

OpenStudy (unklerhaukus):

f(3) = 2f(3 -1) f(3 - 2) = 2f(2) f(1) = 2 x 2 x 1 = ?

OpenStudy (geekfromthefutur):

4 i think

OpenStudy (unklerhaukus):

yeah so that fits the table

OpenStudy (unklerhaukus):

f(4) = 2f(4 -1) f(4 - 2) = 2f(3) f(2) = 2 x 4 x 2 = ?

OpenStudy (geekfromthefutur):

16

OpenStudy (unklerhaukus):

good, lets check then next term f(5) = 2f(5 -1) f(5 - 2) = ?

OpenStudy (geekfromthefutur):

13 i think i may of got that wrong

OpenStudy (geekfromthefutur):

im not sure tht one but for a crazy answer im going to say 128 is the answer

OpenStudy (unklerhaukus):

f(5) = 2f(4) f(3) = ?

OpenStudy (geekfromthefutur):

4096

OpenStudy (geekfromthefutur):

wait no its its 7 right?

OpenStudy (unklerhaukus):

we just found f(4)=16 and f(3) =4

OpenStudy (geekfromthefutur):

oh ok

OpenStudy (geekfromthefutur):

so f(4) = 16 and f(3)= 4 then i would add 16 and 4 together then i would get 20 is tht correct ?

OpenStudy (unklerhaukus):

if we are testing the fourth option f(n) = 2f(n -1) f(n - 2) you have to multiply the terms (not add)

OpenStudy (geekfromthefutur):

omg im srry i meant mutiply not add i would get 64

OpenStudy (unklerhaukus):

yeah f(5) = 2f(4) f(3) = 2 x 16 x 4 = 2 x 64 =

OpenStudy (geekfromthefutur):

128

OpenStudy (unklerhaukus):

is that the same as f(5) of the table?

OpenStudy (geekfromthefutur):

yea

OpenStudy (unklerhaukus):

goody, what about f(6) ?

OpenStudy (geekfromthefutur):

yup its 4096

OpenStudy (unklerhaukus):

so we have checked the last option and all the points fit,

OpenStudy (unklerhaukus):

If we had tried the third option f(1) = 1, f(2) = 2, f(n) = 2f(n -1) + f(n - 2); n > 2 f(3) = 2f(3 -1) + f(3 - 2) = 2 f(2) + f(1) = 2 x 2 + 1 = 5 we see if dosen't match up with the table

OpenStudy (geekfromthefutur):

oh ok so im a little confused tho im looking at my choices of wat the answer could be but its like its not matching

OpenStudy (geekfromthefutur):

oh wow wait never mine it is the last option im srry im a little air headed today :/

OpenStudy (unklerhaukus):

we tried the last option, with f(n) = 2f(n -1) f(n - 2); n > 2 first, and all the points matched the ones on the table

OpenStudy (geekfromthefutur):

ik but i submitted it and it was the was the last one was correct :D:D